40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 190     Entry time: Wed Jul 22 11:09:45 2009
Author: Michelle 
Type: Misc 
Category:  
Subject: Weekly Update 5 

Over the course of the past week I've done a few things. When I began alignment I discovered that our laser was periodically shutting itself down. This was a very perplexing problem for about 2 days before Alberto came in with the diagnosis: the diode was overheating. We'll be sending that in to get it fixed, and we'll also put a heat sink on its casing. Hopefully that will be up and running by the time we really need it.

Right now, Aidan's laser is set up with a 50/50 beamsplitter, and then a half wave plate at each output of that. That way we can run the laser at full power, and each of us can independently adjust the power going into our respective experiments. This is working well while we're aligning things, but it clearly won't work long-term - we need to act directly on the laser's frequency to lock it to a cavity. This may not fare so well for fiber noise suppression.

We have our setup mostly aligned. The beam is going through some steering mirrors, through a lens, into the Faraday Isolator, through another set of steering mirros and a lens, and into the EOM. There is very little loss inside the isolator (putting in ~35 mW and getting out 33-34 mW), however I have not been able to get the power at the output of the EOM higher than ~ 24 mW. I don't think this is normal, but I will check that with people who know better than I do. I think it is probably the fault of poor alignment - the aperture is ~ 2mm and it's about 2 cm from the lens in the middle of the table, so it's hard to reach it or even view it properly to see what needs tweaking.

Over the next week I plan to finish the alignment and hopefully get a lock. I'll get a picture up once the rest of the setup is in place and aligned.

 

Oh, and I also helped clean the lab this week. It's pretty shiny, except for the heaps of garbage boxes now sitting in the hallway. We'll take care of that soon.

ELOG V3.1.3-